OpenStudy (anonymous):

] Part 1: Solve the inequality -3x + 4 less than or greater to x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

hey 1st of all <= stands for less than or "equal to" not greater to

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nedd help or got it?

OpenStudy (anonymous):

Need help solving, \[-3x +4 \le \to x + 8\]

OpenStudy (anonymous):

Whoops I mean -3x + 4 <= x + 8

OpenStudy (anonymous):



OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

I dont need that, I need to solve the inequality

OpenStudy (anonymous):

@dpaInc Can you help please?

OpenStudy (anonymous):

well here it is then\[-3x+4 \le x+8\] \[4-8 \le x+3x\] \[-4 \le 4x .......... -1 \le x\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I dont get the end

OpenStudy (anonymous):

take the 4 on left side so -4/4 = -1 right